OpenStudy (awesomedude2078):

If a car averages 50 1/3 miles per hour, how many hours will it take to go 47 1/2 miles as an improper fraction

OpenStudy (joannablackwelder):

Do you know a relationship between rate, time, and distance?

OpenStudy (awesomedude2078):

can you refresh my memory

OpenStudy (joannablackwelder):

distance = rate times time

OpenStudy (joannablackwelder):

Can you solve that for time?

OpenStudy (awesomedude2078):

time = rate times distance

OpenStudy (sshayer):

\[50\frac{ 1 }{ 3 }=\frac{ 151 }{ 3 }\] \[47\frac{ 1 }{ 2 }=\frac{ 95 }{ 2 }\] \[time~taken ~\to~complete~\frac{ 151 }{ 3 }~miles=1~hour\] \[time~taken~\to~travel ~1~mile=1*\frac{ 3 }{ 151 }~hours.\] \[time~taken~\to~travel \frac{ 95 }{ 2 }~miles=\frac{ 3 }{ 151 } \times \frac{ 95 }{ 2 }=?\]

OpenStudy (awesomedude2078):

is that correct

OpenStudy (joannablackwelder):

 Check your solving step.

OpenStudy (awesomedude2078):

gotta go be right back

OpenStudy (niksbhalla14mar):

Question is simple Speed is 151/3 miles/hr Distance is 95/2 miles Do Time = Distance / Speed = (95/2) / (151/3) = 285/302 Answer is 285/302